OpenStudy (breanneh):

Malia is using an online calculator to calculate the outputs f(n) for different inputs n. The ordered pairs below show Malia's inputs and the corresponding outputs displayed by the calculator: (1, 4), (2, 9), (3, 14), (4, 19) Which of the following functions best represents the rule that the calculator uses to display the outputs? f(n) = 5n − 1 f(n) = 5n + 1 f(n) = 4n + 1 f(n) = 4n − 1

OpenStudy (erak):

So you can just a check yourself by plugging in "n" into the equations using the first number in the given outputs. (n,y)

OpenStudy (mrnood):

the terminology is the same as saying (taking the last example) y=4x-1 looking at the first pair of results : (1,4) this is equivalent to y = 4 when x = 1 so put those values into the equation: 4 = 4(1)-1 = 4-1 =3 You can see that his equation is NOT true - therefore the option d is not the correct answer. Do th esame for the other thre equations and find the one that is true for all the pairs of numbers